How much does it cost to operate a 400w television for 8hours if electrical energy costs php 6.88 per kilowatt
1
Expert's answer
2018-04-02T10:47:08-0400
The total electrical energy used by TV
E=P×t=400 W×8 h=3200 Wh=3.2 kWh
If cost is 6.88 cents per kWh, the amount of money for electricity
3.2×6.88=22 cents